OpenStudy (help_people):

The first-, second-, and third-year enrollment values for a technical school are shown in the table below. Enrollment at a Technical School Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table? The solution to f(x) = s(x) is x = 2,012. The solution to f(x) = s(x) is x = 732. The solution to s(x) = t(x) is x = 2,011. The solution to s(x) = t(x) is x = 710.

OpenStudy (help_people):

A graph of 2 functions is shown below. graph of function f of x equals 2 multiplied by x plus 1 and graph of function g of x equals x cubed plus 2 multiplied by x squared minus x minus 2 Which of the following is an approximate solution for f(x) = g(x)? x = −2 x = 0 x = −1 x = 1

OpenStudy (help_people):

@welshfella

OpenStudy (welshfella):

for the 2nd question i can't be sure how to write it Have you got a picture of the question?

OpenStudy (help_people):

yes :) (a graph ) hold on a sec let me get it

OpenStudy (welshfella):

look at the graph The solution to f(x) = g(x) are where the graph intersect each other There is one of them which is equal to one of the options. Can you pick it out? Look at the 3 intersections and note the values of x.

OpenStudy (help_people):

they intersect at -1 on the x axis?

OpenStudy (help_people):

the graph not they

OpenStudy (welshfella):

yes that's right x = -1 is the answer

OpenStudy (help_people):

it also intersects betweenn -5.5 and -2.75 and 5.5 and 2.75 @welshfella

OpenStudy (welshfella):

its not exactly at x = -1 but very close . The question only asked for an approximate value.

OpenStudy (help_people):

ok :)

OpenStudy (welshfella):

yes but x=-1 is the only one that's in the choices.

OpenStudy (help_people):

ok perfect! could we do the other question now?

OpenStudy (welshfella):

I don't understand it. It talks about functions but gives no details.

OpenStudy (help_people):

it asks what is true about the data ?

OpenStudy (help_people):

in the table are you sure you cannot help ?

OpenStudy (welshfella):

the only thing I can see is for year 2012 f(x) = s(x) ( both are = to 732) so A

OpenStudy (welshfella):

yea thats what it is.

OpenStudy (help_people):

ok i have 2 more and i did them i just want to make sure they are correct?

OpenStudy (welshfella):

sorry I have to go now